# A Survey of Language-Based Communication in Robotics

### v1

@@ 3.3. Robot to Robot Communication
Multi-robot systems often have a communication component, where robots exchange bids, plans, observations, or any other information. Classically this has been a carefully structured process based on predefined protocols. LLMs, however, provide the opportunity to remove these constraints and allow generative models to produce messages in a dialogical manner. From this dialogue emerges a collective intelligence which supports coordination, planning, and knowledge transfer. This approach is particularly appealing as it models a core strength of human groups using the exact same tools; language allows teams of people to organise, robots may be the same. We term this type of system “Robot to Robot Communication”.
@@
Table 3. A selection of works using language for Human-to-Robot tasking. The “LLM” column indicates the most recent LLM they used (not all if they used several variants). Application is listed, as all use language in the same way, but the usecases are often different. Number of agents and whether a Human-in-the-Loop is possible (HITL) are also included.
Modality Paper LLM Application #Agents Human in loop
Role Playing Li et al. LLaMa-7B Various inc. code, maths, science 2
Chan et al. GPT-4 Debate to give advice 3+
Hong et al. GPT-4 Software development team 2+
Wang et al. GPT-4 Finding objects in ALFWorld 2
Fu et al. GPT-4 & Claude Bartering (buyer/seller) 2
Nascimento et al. GPT-4 Bartering (buyer/seller) 2+
Zhang et al. GPT-3.5 Various inc. quiz, maths, chess 2+
Zeng et al. Various Various captioning tasks 3  [^3] The authors include people as a means of input to their language model component
Wu et al. GPT-4 Various (mostly game theory) 2+
Inter-Agent Coordination Park et al. GPT-3.5 Model a community 2+
Bigazzi et al. CLIP Navigation and captioning 2+
Zhang et al. GPT-4 Fetching items in a home 2
Wu et al. GPT-4 Various inc. maths, code, chess 2+
Chen et al. GPT-4 Complex plans as advice, coding, etc. 1+
Papangelis et al. N/A Dialogues 2
Du et al. N/A Reasoning 2
Inter-Robot Kannan et al. GPT-4 Planning and task allocation 2+
Mandi et al. GPT-4 Planning and coordination 2+
Hunt et al. GPT-4 Planning and coordination 2+
3.3.1. Role Playing
@@
Language models such as GPT generally have a “system message”. This is a prompt which the model considers with each message it sends; it can be thought of as the agent’s “ego” (OpenAI, 2023). For example, in assistant-style systems like ChatGPT, this is often set to something like “you are a helpful assistant” (Chi et al., 2023). In practise these can be very long and complex and are often not fully published, but overall the system prompt defines the high-level behaviour and themes of the LLM. This can also be used to define a character for the agent, such as a tendency to be more direct, humorous, or polite. More formally, recognised personality types can be modelled and reflected by the LLM’s outputs (Safdari et al., 2023). This practise is called “role playing” and allows a human to interact with a specialised agent, or can be used to create a dialogue between two chatbots, each of whom acts as though talking to a human, whilst assuming their defined character or style (Li et al., 2023b). An approach of this form can allow an LLM to critique itself for iterative improvement (Ye et al., 2023), or can use dialogue to encourage one agent to instruct another step-by-step on how to solve a task (Wang et al., 2023b). This can be expanded beyond two agents to a team, each of whom is given a slightly different role to play (Chan et al., 2023). This mimics the way human teams are often constructed with a diverse combination of skillsets, backgrounds, and personality types (Pieterse et al., 2006; Pollock, 2009), which is often suggested to be beneficial in certain settings (Higgs et al., 2005; Meslec and Curşeu, 2015; Pieterse et al., 2006). Although typically decentralised and peer-to-peer, such teams can also be modelled after hierarchical structures. For example, a team of agents each with a different, role-playing specialism can be built, and as each works on a group solution they alter it and pass it on to their subordinate. An engineer can write code and pass it to a QA tester who validates its effectiveness (Hong et al., 2023; Qian et al., 2023). This allows different specialisms to be combined, although it is worth considering that the system prompts do not alter an agent’s actual knowledge or skills. Future systems may finetune different models to this end. These processes often require substantial human guidance, although a teacher-student system can be used to automate some of this process (Rasal, 2024). Similarly, feedback and criticism from experience can be directly used to update such agents, allowing a group to criticise each other and collectively improve (Fu et al., 2023), or else to compete with each other and each attempt to get the best outcome from a collective exchange (Nascimento et al., 2023). The process of forming optimal teams of differently configured language agents is being actively explored and may become a crucial part of coalition formation in future (Zhang et al., 2023g). Finally, multiple ways of interpreting the environment can be combined and treated as separate parts of the total model. This allows other types of data such as images and audio to be directly factored into the collective model. Dialogue can be an effective interface between these models as they are each fundamentally text-orientated (Zeng et al., 2022)
3.3.2. Inter-Agent Coordination
@@
Most closely matching an ad-hoc human group, inter-agent coordination describes groups of agents who share a common central repository such as a message board. Any agent can view this board, consider the current messages, and write a reply which is posted. This model arises from ChatGPT, which models an agent in dialogue with a human. By restructuring the conversation so that each agent perceives other agents as aspects of its human interlocutor, the conversation can function naturally, with each agent participating and considering the suggestions of others. This has a key advantage of being feasible to organise with heterogeneous agents and no prior setup (provided a chat-board can be maintained, the only constraint is that each agent can read and post). Additionally, conversation forms the ideal basis for the merging and sharing of diverse knowledge and experience sets. Agents may have experienced different things and have built up a different understanding of their environment. It is not always practical to merge models, exchange maps, or upload images. In conversation, however, agents can condense this information into language which encapsulates much of that knowledge and insert it into the group consciousness.
@@
A foundational work of this form places a set of agents in a simulated environment where they can move, act, and engage in dynamic dialogues with other agents. This forms a community in which each agent acts according to its identity and experiences; knowledge such as current events and gossip diffuse throughout the group. Although this work has no direct utility, it acts as a robust proof-of-concept for language-based agent collectives (Park et al., 2023). One challenge for this approach is deciding when to speak; language models are slow to run, and conversation requires efficiency. Agents should not constantly flood the network with messages and should instead determine whether they have something worth saying, otherwise listening to others. This can be achieved with a separate, lower-weight model which determines if the main language generating model should be invoked (Bigazzi et al., 2023). Another consideration from a work that does not employ LLMs is to direct messages to other agents as questions, prompting them to update the sender with what they know (Papangelis et al., 2019). As language is a dense storage of meaning, it can be useful in scenarios where decentralised agents have a high cost for communication. A single sentence can infer a large amount of knowledge about environment, actions, and intention to other agents (Zhang et al., 2023c). Such approaches can be extended to allow agents to internally reason and predict their neighbour’s actions and intentions. This can be done with direct language exchange, or can be used to consider one’s neighbour using an LLM without explicitly communicating with them; essentially using language to facilitate robot-to-robot coordination and teamwork (Zhang et al., 2023h). Different architectures can be explored for inter-agent coordination, such as a hierarchy with a supervisor agent or a decentralised joint chat. Such architectures also facilitate the placement of a human in any position, essentially extending this approach to encompass human-robot and robot-human in the same system (Wu et al., 2023b). The same fundamental approach can be used for the entire structure of the team while it pursues its goal. A task can be provided to agents who recruit the best team of experts, decide on their strategy, and enact the plan; language is the basis of each step (Chen et al., 2023e). The core intuition here is that a team of communicating agents can reach a better solution through debate than a single agent can from thinking alone (Du et al., 2023a)
3.3.3. Inter-Robot Coordination
@@
The coordination of robots using language is not a well explored area, but a few works do exist. The most clear example is RoCo, which sets a team of robot arms simple tasks and has them coordinate using language. Each robot speaks to the others in a shared chat log to propose and approve each others plans. It also responds to mistakes such as collisions (Mandi et al., 2023). Another work has two robots discuss their solution to a high-level task in debate format, resulting in paths of rooms in an environment. This approach also allows human approval of proposed plans by calling the human once agents agree; the human can also intervene during execution (Hunt et al., 2024).
3.4. Robot Control and Reasoning
@@
Language and language-based models can also be used internally withing a single robot. This can take a variety of forms, but in general these approaches seek to utilise the intelligence of LLMs within the robot’s own control and reasoning processes.
@@
Table 4. A selection of works using language for Human-to-Robot tasking. The “LLM” column indicates the most recent LLM they used (not all if they used several variants). Italics indicate a simulator, and “language used to” is left high level to allow papers to be further subgrouped.
Modality Paper LLM Robot/Sim Language used to
Transformer Robotics Brohan et al. RT-2 7-DoF Robot Arm Describe task
Driess et al. PALM-E Various Describe and plan
Padalkar et al. RT-X 7-DoF Robot Arm Describe task
Ahn et al. AutoRT 7-DoF Robot Arm Describe task & alignment
Language Architecture Wang et al. GPT-4 Various Acts on joints
Huang et al. InstructGPT & CLIPort UR5e Arm Plan and act
Ulmer et al. OpenLlama N/A Guide through behaviours
Lykov et al. Alpaca 7B N/A Build behaviour trees
Cao and Lee GPT-3.5 N/A Build behaviour trees
Wei et al. GPT-4 & PaLM N/A Planning and reasoning
Yao et al. GPT-4 N/A Planning and reasoning
Du et al. PaLM-E Various Planning and acting
Besta et al. GPT-3.5 N/A Planning and reasoning
Xie et al. GPT-4 Unitree Go1 Planning and reasoning
Gu et al. GPT-4 Clearpath Jackal Planning and reasoning
Dynamic Adaptation Shirai et al. GPT-4 Robot Arm Re-prompting from error messages
Shek et al. GPT-4 spot-mini-mini Modify terrain type
Skreta et al. GPT-3 Franka Panda Reprompting from error messages
Jin et al. GPT-4 Franka Research 3 Environment feedback to LLM
Bhat et al. GPT-4 & PALM-2 Franka Research 3 High and Low-level planning
Yao et al. GPT-3 N/A Q-A
3.4.1. Transformer Robotics
@@
Language models are typically built on the transformer architecture, which is defined by an attention mechanism and connections between the encoder and decoder. Although not exclusive to language, generally they are used to map language inputs to language outputs (Vaswani et al., 2017). A similar architecture can however be modified by making the output space a language-based representation of robot parameters (Brohan et al., 2022). Combining this with a Vision-Language Model creates an architecture which takes in an image of the environment and a textual task, and maps this in one step to a robot action which can be executed in real life (Brohan et al., 2023a). This can generalise to different tasks using the same policy, and can also generalise to different robots, essentially transferring learning across platforms and tasks (Padalkar et al., 2023). Further to this, language-based policies can be combined with pre-written skills and human teleoperation, each of which exist as part of a total architecture for training and execution (Ahn et al., 2024). These approaches offer an opportunity for language to bridge the gap between multiple domains to form a generalised, platform-agnostic, multimodal model for robotic systems (Driess et al., 2023).
3.4.2. Language-based Architectures
@@
Some architectures use language internally or to facilitate other means of robotic control. This can be highly effective in bringing together different component models and combining them through a language interface. An example is InnerMonologue (Huang et al., 2022b), which uses language to create an internal conversation between different aspects of the robot. The robot might call the scene descriptor to announce what it can see, after which the robot checks with the human what they want, and then the robot executes an action and asks the success detector if it achieved its goal. Each part of the architecture is “talking” to each other within the overall structure. Such “self talk” is a powerful tool and may be useful in a number of domains including navigation (Xie et al., 2023). Other approaches use language as the entire policy by providing textual descriptions of the environment and actions, and use the LLM to generate joint parameters in one step (Wang et al., 2023d). More broadly, some systems create a complex architecture with each component being language-based. From environment representation to action, language is the representation space (Wang et al., 2024a). Another way of doing this is to pre-generate responses to your expectations, and use the LLM to best fit the situation to those responses. For example, a dialogue tree can be created and the LLM is asked to match a user’s input to whichever response most closely matches the expected answers on the tree (Ulmer et al., 2024), or alternatively the most likely user inputs can be predicted and responses can be pre-generated; something that could feasibly done for robotics as most user requests are likely to be predictable in the context of the overall mission or goal (Chi et al., 2023). Some approaches use LLMs in conjunction with another representation of information such as using graphs to perform basic mathematics (Cao, 2023), or representing knowledge in a graph structure to be queried and reasoned through as a tree (Li et al., 2024).
@@
LLMs can also be employed to process only the abstract and high-level parts of the problem, for example by deciding where to go or guessing where an object is likely to be, with all other actions performed using more standard approaches (Gu et al., 2023b). An alternative method is to use language to generate a policy that is then used. The classic method of doing this is to generate a behaviour tree which then controls the robot (Cao and Lee, 2023; Lykov and Tsetserukou, 2023). The behaviour tree approach can also be used to control a team of robots (Lykov et al., 2023). Being next-symbol predictors, LLMs sometimes struggle to reason about questions, especially when careful consideration of several factors is involved. A proposed solution is to encourage them to give their reasoning as they answer, a “chain-of-thought” approach (Wei et al., 2022). An extension of this approach is to create a “tree-of-thoughts” which considers multiple paths and includes backtracking to find the best reasoning (Yao et al., 2024b). An improvement on this tree structure is a graph structure which also supports backtracking and aggregation of thoughts into the final plan (Besta et al., 2023). This can further be developed by combining visual information; by breaking down the task into a tree structure and employing text-to-video models, a complete plan can be generated including both task specifications and video demonstrations extrapolated from the initial state and future actions (Du et al., 2023b). The trend of equipping LLMs with understanding of other data structures is common across multiple works and appears to be a core means of empowering language-based systems (Wang et al., 2024d)
@@
These complex decision making problems, especially within an environment which is difficult to interpret and understand make video games a natural test case for these approaches. Language has been applied as a planning tool in Minecraft (Zhu et al., 2023), and as an instruction space for planning and action for agents whose learning can transfer across multiple games (Raad et al., 2024). Each of these developments increases the complexity of language-based environmental reasoning that can be done, a problem that is discussed later in section 5.2.1.
3.4.3. Dynamic Adaptation with Language

### v3

@@ 3.3 Robot to Robot Communication
Multi-robot systems often have a communication component, where robots exchange bids, plans, observations, or any other information. Classically this has been a carefully structured process based on predefined protocols. LLMs, however, provide the opportunity to remove these constraints and allow generative models to produce messages in a dialogical manner. From this dialogue emerges a collective intelligence which supports coordination, planning, and knowledge transfer. This approach is particularly appealing as it models a core strength of human groups using the exact same tools; language allows teams of people to organise, and the same may be true for robots. We term this type of system “Robot to Robot Communication”.
@@
Table 3: A selection of works using language for Robot-to-Robot communication. The “LLM” column indicates the most recent LLM they used (not all if they used several variants). Application is listed, as all use language in the same way, but the use cases are often different. The number of agents and whether a Human-in-the-Loop is possible (HITL) are also included.
Modality Paper LLM Application #Agents Human in loop Role Playing Li et al., 2023a LLaMa-7B Reasoning 2 Chan et al., (2023) GPT-4 Debating 3+ Hong et al., (2023) GPT-4 Developing Software 2+ Wang et al., 2023c GPT-4 Finding objects in ALFWorld 2 Fu et al., (2023) GPT-4 & Claude Bartering (buyer/seller) 2 Nascimento et al., (2023) GPT-4 Bartering (buyer/seller) 2+ Zhang et al., 2023d GPT-3.5 Reasoning 2+ Zeng et al., (2022) Various Captioning 3  [^3] The authors include people as a means of input to their language model component Wu et al., 2024c GPT-4 Planning 2+ Inter-Agent Coordination Park et al., (2023) GPT-3.5 Modelling a community 2+ Bigazzi et al., (2023) CLIP Navigating and captioning 2+ Zhang et al., 2023c GPT-4 Fetching items in a home 2 Wu et al., 2023b GPT-4 Reasoning 2+ Chen et al., 2023c GPT-4 Reasoning 1+ Papangelis et al., (2019) N/A Planning through dialogue 2 Du et al., 2023a N/A Reasoning 2 Inter-Robot Kannan et al., (2023) GPT-4 Planning and allocating tasks 2+ Mandi et al., (2023) GPT-4 Planning and coordinating 2+ Hunt et al., (2024) GPT-4 Planning and coordinating 2+
3.3.1 Role Playing
@@
Language models such as GPT generally have a “system message”. This is a prompt that the model considers with each message it sends; it can be thought of as the agent’s “ego” OpenAI, (2023). For example, in assistant-style systems like ChatGPT, this is often set to something like “you are a helpful assistant” Chi et al., (2023). In practise these can be very long and complex and are often not fully published, but overall the system prompt defines the high-level behaviour and themes of the LLM. This can also be used to define a character for the agent, such as a tendency to be more direct, humorous, or polite. More formally, recognised personality types can be modelled and reflected by the LLM’s outputs Safdari et al., (2023). This practice is called “role playing” and allows a human to interact with a specialised agent, or can be used to create a dialogue between two chatbots, each of whom acts as though talking to a human, whilst assuming their defined character or style Li et al., 2023a . An approach of this form can allow an LLM to critique itself for iterative improvement Ye et al., (2023), or can use dialogue to encourage one agent to instruct another step-by-step on how to solve a task Wang et al., 2023c . This can be expanded beyond two agents to a team, each of whom is given a slightly different role to play Chan et al., (2023). This mimics the way human teams are often constructed with a diverse combination of skillsets, backgrounds, and personality types Pieterse et al., (2006); Pollock, (2009), which is often suggested to be beneficial in certain settings Higgs et al., (2005); Meslec and Curșeu, (2015). Although typically decentralised and peer-to-peer, such teams can also be modelled after hierarchical structures. A team of agents each with a different, role-playing specialism can be built, and as each works on a group solution they alter it and pass it on to their subordinate. For example an engineer can write code and pass it to a QA tester who validates its effectiveness, and so on Hong et al., (2023); Qian et al., (2023). This allows different specialisms to be combined, although it is worth considering that the system prompts do not alter an agent’s actual knowledge or skills. Future systems may finetune different models to this end. These processes often require substantial human guidance, although a teacher-student system can be used to automate some of this process Rasal, (2024). Similarly, feedback and criticism from experience can be directly used to update such agents, allowing a group to criticise each other and collectively improve Fu et al., (2023), or else to compete with each other and each attempt to get the best outcome from a collective exchange Nascimento et al., (2023). The process of forming optimal teams of differently configured language agents is being actively explored and may become a crucial part of coalition formation in future Zhang et al., 2023d . Finally, multiple ways of interpreting the environment can be combined and treated as separate parts of the total model. This allows other types of data such as images and audio to be directly factored into the collective model. Dialogue can be an effective interface between these models as they are each fundamentally text-orientated Zeng et al., (2022).
3.3.2 Inter-Agent Coordination
@@
Most closely matching an ad-hoc human group, inter-agent coordination describes groups of agents who share a common central repository such as a message board. Any agent can view this board, consider the current messages, and write a reply which is posted. This concept arises from ChatGPT, which models an agent conversing with a human. By restructuring the conversation so that each agent perceives other agents as aspects of its human interlocutor, the conversation can function naturally, with each agent participating and considering the suggestions of others. This has a key advantage of being feasible to organise with heterogeneous agents and no prior setup (provided a chat-board can be maintained, the only constraint is that each agent can read and post). Additionally, conversation forms the ideal basis for the merging and sharing of diverse knowledge and experience. Agents may have experienced different things and have built up a different understanding of their environment. It is not always practical to merge models, exchange maps, or upload images. In conversation, however, agents can condense this information into language which encapsulates much of that knowledge, and insert it into the group consciousness.
@@
A foundational work of this form places a set of agents in a simulated environment where they can move, act, and engage in dynamic dialogues with other agents. This forms a community in which each agent acts according to its identity and experiences; knowledge such as current events and gossip diffuses throughout the group. Although this work has no direct utility, it acts as a robust proof-of-concept for language-based agent collectives Park et al., (2023). One challenge for this approach is deciding when to speak; language models are slow to run, and conversation requires efficiency. Agents should not constantly flood the network with messages and should instead determine whether they have something worth saying, otherwise listening to others. This can be achieved with a separate, lower-weight model which determines if the main language-generating model should be invoked Bigazzi et al., (2023). Another consideration from a work that does not employ LLMs is to direct messages to other agents as questions, prompting them to update the sender with what they know Papangelis et al., (2019). As language is a dense storage of meaning, it can be useful in scenarios where decentralised agents have a high cost for communication. A single sentence can infer a large amount of knowledge about the environment, actions, and intentions to other agents Zhang et al., 2023c . Such approaches can be extended to allow agents to internally reason and predict their neighbours’ actions and intentions. This can be done with direct language exchange, or can be used to consider one’s neighbour using an LLM without explicitly communicating with them; essentially using language to facilitate robot-to-robot coordination and teamwork Zhang et al., 2023b . Different architectures can be explored for inter-agent coordination, such as a hierarchy with a supervisor agent or a decentralised joint chat. Such architectures also facilitate the placement of a human in any position, essentially extending this approach to encompass the human-robot and robot-human categories in the same system Wu et al., 2023b . The same fundamental approach can be used for the entire structure of the team while it pursues its goal. A task can be provided to agents who recruit the best team of experts, decide on their strategy, and enact the plan; language is the basis of each step Chen et al., 2023c . The core intuition here is that a team of communicating agents can reach a better solution through debate than a single agent can from thinking alone Du et al., 2023a .
3.3.3 Inter-Robot Coordination
@@
The coordination of robots using language is not a well-explored area, but a few works do exist. The most clear example is RoCo, which sets a team of robot arms simple tasks and has them coordinate using language. Each robot speaks to the others in a shared chat log to propose actions and approve each others’ plans. It also responds to mistakes such as collisions Mandi et al., (2023). Another work has two robots discuss their solution to a high-level task in a debate format, resulting in paths of rooms through an environment. This approach also allows human approval of proposed plans by calling the human once agents agree; the human can also intervene during execution Hunt et al., (2024).
3.4 Robot Control and Reasoning
@@
Language and language-based models can also be used internally within a single robot. This can take a variety of forms, but in general, these approaches seek to utilise the intelligence of LLMs within the robot’s own control and reasoning processes.
@@
Table 4: A selection of works using language for robot control and reasoning. The “LLM” column indicates the most recent LLM they used (not all if they used several variants). Italics indicate a simulator, and “Application” is left high level to allow papers to be further subgrouped.
Modality Paper LLM Physical Robot / Simulation Application Transformer Robotics Brohan et al., (2022) RT-1 7-DoF Robot Arm Describing task Brohan et al., 2023a RT-2 7-DoF Robot Arm Describing task Driess et al., (2023) PALM-E Various Describing and planning Padalkar et al., (2023) RT-X 7-DoF Robot Arm Describing task Ahn et al., (2024) AutoRT 7-DoF Robot Arm Describing task & alignment Language Architecture Wang et al., 2023f GPT-4 Various Acting on joints Huang et al., 2022b InstructGPT & CLIPort UR5e Arm Planning and acting Ulmer et al., (2024) OpenLlama N/A Guiding through behaviours Lykov et al., (2023) Alpaca 7B N/A Building behaviour trees Cao and Lee, (2023) GPT-3.5 N/A Building behaviour trees Wei et al., (2022) GPT-4 & PaLM N/A Planning and reasoning Yao et al., 2024a GPT-4 N/A Planning and reasoning Du et al., 2023b PaLM-E Various Planning and acting Besta et al., (2023) GPT-3.5 N/A Planning and reasoning Xie et al., (2023) GPT-4 Unitree Go1 Planning and reasoning Gu et al., 2023b GPT-4 Clearpath Jackal Planning and reasoning Dynamic Adaptation Shirai et al., (2023) GPT-4 Robot Arm Re-prompting from error messages Shek et al., (2023) GPT-4 spot-mini-mini Modifying terrain type Skreta et al., (2023) GPT-3 Franka Panda Reprompting from error messages Jin et al., (2023) GPT-4 Franka Research 3 Providing feedback Bhat et al., (2024) GPT-4 & PALM-2 Franka Research 3 Planning Yao et al., (2023) GPT-3 N/A Question-Answering
3.4.1 Transformer Robotics
@@
Language models are typically built on the transformer architecture, which is defined by an attention mechanism and connections between the encoder and decoder. Although not exclusive to language, generally they are used to map language inputs to language outputs Vaswani et al., (2017). A similar architecture can, however, be modified by making the output space a language-based representation of robot parameters Brohan et al., (2022). Combining this with a Vision-Language Model (VLM) creates an architecture which takes in an image of the environment and a textual task, and maps this in one step to a robot action which can be executed in real life Brohan et al., 2023a . This can generalise to different tasks using the same policy, and can also generalise to different robots, essentially transferring learning across platforms and tasks Padalkar et al., (2023). Further to this, language-based policies can be combined with pre-written skills and human teleoperation, each of which exists as part of a total architecture for training and execution Ahn et al., (2024). These approaches offer an opportunity for language to bridge the gap between multiple domains to form a generalised, platform-agnostic, multimodal model for robotic systems Driess et al., (2023).
3.4.2 Language-based Architectures
@@
Some architectures use language internally or to facilitate other means of robotic control. This can be highly effective in bringing together different component models and combining them through a language interface. An example is InnerMonologue Huang et al., 2022b , which uses language to create an internal conversation between different aspects of the robot. The robot might call the scene descriptor to announce what it can see, after which the robot checks with the human what they want, and then the robot executes an action and asks the success detector if it has achieved its goal. Each part of the architecture is “talking” to each other within the overall structure. Such “self-talk” is a powerful tool and may be useful in a number of domains including navigation Xie et al., (2023). Other approaches use language as the entire policy by providing textual descriptions of the environment and actions, and use the LLM to generate joint parameters in one step Wang et al., 2023f . More broadly, some systems create a complex architecture with each component being language-based. From environment representation to action, language is the representation space Wang et al., 2024c . Another way of doing this is to pre-generate responses to your expectations and use the LLM to best fit the situation to those responses. For example, a dialogue tree can be created and the LLM is asked to match a user’s input to whichever response most closely matches the expected answers on the tree Ulmer et al., (2024). Alternatively, the most likely user inputs can be predicted and responses can be pre-generated; this is something that could feasibly done for robotics as most user requests are likely to be predictable in the context of the overall mission or goal Chi et al., (2023). Some approaches use LLMs in conjunction with another representation of information such as using graphs to perform basic mathematics Cao, (2023), or representing knowledge in a graph structure to be queried and reasoned through as a tree Li et al., (2024).
@@
LLMs can also be employed to process only the abstract and high-level parts of the problem, for example by deciding where to go or guessing where an object is likely to be, with all other actions performed using more standard approaches Gu et al., 2023b . An alternative method is to use language to generate a policy that is then used. The classic method of doing this is to generate a behaviour tree which then controls the robot Cao and Lee, (2023); Lykov and Tsetserukou, (2023). The behaviour tree approach can also be used to control a team of robots Lykov et al., (2023). Being next-symbol predictors, LLMs sometimes struggle to reason about questions, especially when careful consideration of several factors is involved. A proposed solution is to encourage them to give their reasoning as they answer — a “chain-of-thought” approach Wei et al., (2022). An extension of this approach is to create a “tree-of-thoughts” which considers multiple paths and includes backtracking to find the best reasoning Yao et al., 2024a . An improvement on this tree structure is a graph structure which also supports backtracking and aggregation of thoughts into the final plan Besta et al., (2023). This can further be developed by combining visual information; by breaking down the task into a tree structure and employing text-to-video models, a complete plan can then be generated including both task specifications and video demonstrations extrapolated from the initial state and future actions Du et al., 2023b . The trend of equipping LLMs with an understanding of other data structures is common across multiple works and appears to be a core means of empowering language-based systems Wang et al., 2024d .
@@
These complex decision-making problems, especially within an environment which is difficult to interpret and understand make video games a natural test case for these approaches. Language has been applied as a planning tool in Minecraft Zhu et al., (2023), and as an instruction space for planning and action for agents whose learning can transfer across multiple games Raad et al., (2024). Each of these developments increases the complexity of language-based environmental reasoning that can be performed, a problem that is discussed later in section 5.2.1.
3.4.3 Dynamic Adaptation with Language